OpenStudy (anonymous):

Perform the modular arithmetic operation. 8 ∙ 8 (mod 6) =

Parth (parthkohli):

\[8 \quad(\text{mod }6) \implies \text{Remainder in } 8\div 6 \]

Parth (parthkohli):

Then you multiply 8 to that.

Parth (parthkohli):

Recall remainders? The basic-est concept I learnt in elementary Mathematics.

Parth (parthkohli):

Maybe it's something else. I feel that your question is\[(8 \cdot 8)\mod6 \]

OpenStudy (anonymous):

i got 1.33333 for 8/6 .. is this right?

OpenStudy (anonymous):

that is my question..?

Parth (parthkohli):

nvm... Your question is \(8\cdot8\mod{6}\)

Parth (parthkohli):

You first find \(8 \cdot 8\). That'd be the same as \(8 \times 8\)

OpenStudy (anonymous):

64

Parth (parthkohli):

Then perform the long division method to find \(64 \div 6\). Use draw tool...

OpenStudy (anonymous):

soooo i got 10.66667 ?

Parth (parthkohli):

No! Use long division!